OpenStudy (eli_moses):

What is molality of glucose having aqueous solution density is 1 gm/ml and 2 molal?

OpenStudy (eli_moses):

question by @akshay1234

OpenStudy (anonymous):

@xoalybabyxo

OpenStudy (anonymous):

i also tried but i am not sure about my ans

OpenStudy (anonymous):

i will give you medal and be fan if you do this

OpenStudy (anonymous):

@alphadxg @sweetburger @Somy @Brainybeauty

OpenStudy (eli_moses):

@Hero @AccessDenied @thomaster

OpenStudy (somy):

i suppose 2 is mole?

OpenStudy (anonymous):

no it is molal means molality

OpenStudy (somy):

then what is the q? if molality is given what do u need to find?

OpenStudy (anonymous):

we have to find molarity

OpenStudy (somy):

ohhhh

OpenStudy (anonymous):

this are concentration unit

OpenStudy (somy):

alright molarity= mole/volume density= mass/volume volume= mass/density molality= mole/mass mole= molality*mass thus we can make a formula \[molarity= \frac{ molality \times mass }{ mass/density }\] so \[\frac{ molality \times mass }{ 1} \div \frac{ mass }{ density}\] so \[\frac{ molality \times mass }{ 1} \times \frac{ density }{ mass}\] so masses get cancelled and we are left with \[molarity= molality \times density \]

OpenStudy (somy):

bit of playing with formula and here u go a new formula lol

OpenStudy (somy):

do u get it? @Eli_Moses

OpenStudy (somy):

@akshay1234

OpenStudy (eli_moses):

i get it

OpenStudy (somy):

good :))

OpenStudy (eli_moses):

thanks

OpenStudy (somy):

you are welcome :))